Differentiate with respect to
x: f(x)=(ax 3 +bx)
1
Expert's answer
2014-04-23T06:22:20-0400
Using the table of derivatives and the rule that the derivative of the sum equals the sum of derivatives we get
f'(x)=(ax^3+bx)'=(ax^3)'+(bx)'=3ax^2+b